OpenStudy (anonymous):

Solve for x: the quantity of x plus 4 all over 2 = 7

OpenStudy (anonymous):

(x+4)/2 = 7

OpenStudy (anonymous):

(x+4)/2 =7 x+4 = 14 x = 10

OpenStudy (anonymous):

Yes^

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

give me a medal.. please:)

OpenStudy (anonymous):

nvm!

OpenStudy (anonymous):

thx!

OpenStudy (anonymous):

i just gave a medal:)

OpenStudy (anonymous):

i no! i saw that right after i asked:) lol